# Implicit Neural Representations for Robust Joint Sparse-View CT Reconstruction

### v1

@@ Abstract
Computed Tomography (CT) is pivotal in industrial quality control and medical diagnostics. Sparse-view CT, offering reduced ionizing radiation, faces challenges due to its under-sampled nature, leading to ill-posed reconstruction problems. Recent advancements in Implicit Neural Representations (INRs) have shown promise in addressing sparse-view CT reconstruction. Recognizing that CT often involves scanning similar subjects, we propose a novel approach to improve reconstruction quality through joint reconstruction of multiple objects using INRs. This approach can potentially leverage both the strengths of INRs and the statistical regularities across multiple objects. While current INR joint reconstruction techniques primarily focus on accelerating convergence via meta-initialization, they are not specifically tailored to enhance reconstruction quality. To address this gap, we introduce a novel INR-based Bayesian framework integrating latent variables to capture the inter-object relationships. These variables serve as a dynamic reference throughout the optimization, thereby enhancing individual reconstruction fidelity. Our extensive experiments, which assess various key factors such as reconstruction quality, resistance to overfitting, and generalizability, demonstrate significant improvements over baselines in common numerical metrics. This underscores a notable advancement in CT reconstruction methods.
@@ 1 Introduction
Computed Tomography (CT) is a crucial non-invasive imaging technique, extensively employed in medical diagnostics and industrial quality control. In CT, an object’s internal structure is reconstructed from X-ray projections captured at multiple angles, posing a complex inverse problem. In certain situations, reducing the number of CT measurements can yield advantages such as decreased radiation exposure and enhanced production throughput. However, this sparsity in measurements complicates the reconstruction process, making it an ill-posed inverse problem. Such challenges arise not only in CT reconstruction but also across diverse computational tasks. Therefore, while our study centers on sparse-view CT reconstruction, the core ideas are transferable to numerous inverse problems.
@@
Different strategies have been developed to address the challenges of sparse-view CT reconstruction by incorporating auxiliary information. Supervised learning techniques learn mappings from measurements to images (Zhang et al., 2018; Han & Ye, 2018; Zhu et al., 2018; Wu et al., 2021). These methods train on extensive datasets containing dense measurements, synthesizing the sparse view through physical models of the measurement process. Alternatively, Song et al. (2022) introduces an approach that directly learns the image distribution from high-quality reconstructed images using generative models. However, their method still relies on a large, domain-specific dataset, that closely matches the target of reconstruction. Dependence on large datasets introduces practical limitations to supervised learning techniques and generative methods, particularly when collecting large-scale and high-quality data is intractable, such as due to privacy regulation in medical cases or timeliness in industrial settings. There are also some studies leverage heuristic image priors, e.g. Total Variation (TV) (Sidky & Pan, 2008; Liu et al., 2013; Zang et al., 2018). While effective, these methods lack the ability to incorporate domain-specific enhancement. Another approach involves using a few dense-view images as priors for reconstruction (Chen et al., 2008; Shen et al., 2022). However, this method hinges on the availability of dense-view images similar to the target object, which is often impractical when the appearance of the object to be reconstructed is unknown.
On a different tangent, many works explore the potential of implicit neural representations (INRs). Thanks to the continuous representation nature of INRs, these methods have consistently delivered promising reconstruction results with limited number of measurements (Zang et al., 2021; Zha et al., 2022; Rückert et al., 2022; Wu et al., 2023b). Given INRs’ proven capabilities in CT reconstruction and the known advantages of leveraging auxiliary information, we seek to merge these two paradigms. Aiming for broad applicability, we avoid reliance on pre-curated datasets. Instead, we draw inspiration from the fact that modern CT machines routinely scan similar subjects, such as patients in hospitals or analogous industrial products. This commonality drives us to explore a novel question in this research:
@@
Figure 1 contrasts the methodologies of individual reconstruction against joint reconstruction. In our exploration of this research avenue, we found that several existing methods can be adapted for our purpose (Zhang et al., 2013; Ye et al., 2019; Tancik et al., 2021; Martin-Brualla et al., 2021; Kundu et al., 2022). Although these previous works have exploited the statistical regularities among different objects borne in the INR networks’ weights, they often target different problems such as convergence rate (Tancik et al., 2021; Kundu et al., 2022). In terms of improving reconstruction quality, these methods cannot fully capitalize on the available statistical regularities, as shown in experiments Section 5.
@@
Our Contributions: i) We explore a novel problem of INR-based joint reconstruction in the context of CT imaging, supported by a comprehensive review of existing methods that could be adapted to address this challenge. ii) We propose a principled Bayesian framework that adaptively integrates prior information throughout the training process for enhanced reconstruction quality. iii) Through extensive experiments, we evaluate various facets of reconstruction performance using common numerical metrics. Our results establish that our method significantly outperforms the compared INR-based baselines, suggesting notable advancements in the field of CT reconstruction.
@@ 2 Related Work
We briefly outline key studies related to our focal areas, with a comprehensive understanding of INR and Neural Radiance Fields (NeRF) available in the survey (Tewari et al., 2022).
Coordinates-based Multi-Layer Perceptrons (MLPs) have transitioned from traditional discrete representations to INRs by addressing high-frequency function detailing issues (Jacot et al., 2018; Tancik et al., 2020). NeRF, a state-of-the-art INR approach, models continuous scenes using spatial coordinates and viewing angles, incorporating transmittance effects during ray-tracing (Mildenhall et al., 2021; Barron et al., 2021; 2022). Building on the groundwork, several techniques have emerged to leverage the shared information across different objects or scenarios. NeRF-wild (Martin-Brualla et al., 2021) differentiated between static and transient scene aspects, an approach echoed in video representations (Li et al., 2021; Mai & Liu, 2022). Tancik et al. (2021) introduced meta-learning techniques to learn a meta initialization which converges faster than training from scratch. They have also applied their method to CT reconstruction. For the same purpose of accelerating convergence, Kundu et al. (2022) introduced federated learning to obtain the meta initialization. Chen & Wang (2022) proposed using transformers as the meta-learners.
@@ 3 Problem Statement and Preliminaries
Mathematically, the CT acquisition process can be formulated as a linear equation: $\bm{y}=\bm{A}\bm{x}+\bm{\epsilon}$, where $\bm{x}\in{\mathbb{R}}^{m}$ represents the unknown object of interest and $\bm{y}\in{\mathbb{R}}^{n}$ symbolizes the noisy measurements. These measurements arise from the interaction between the measurement matrix $\bm{A}\in{\mathbb{R}}^{n\times m}$ and the object, where $\bm{\epsilon}\in{\mathbb{R}}^{n}$ accounts for the associated measurement noise. The task in CT is to infer the unknown object $\bm{x}$ from the acquired CT measurements $\bm{y}$. The inherent challenge lies in the common sparsity of these measurements, resulting in $m>n$. This makes the reconstruction problem ill-posed. Meanwhile, measurement noise can further corrupt the reconstruction results.
The INR designed for CT reconstruction is a function $f_{\bm{w}}:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}^{1}$ parameterized by $\bm{w}$. It maps the spatial coordinates of the object to its intensity in a continuous three-dimensional space. INR consists of two components, formulated as $f_{\bm{w}}=\mathcal{M}\circ\gamma$. Here, $\gamma:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}^{d}$ servers as the position encoding (Tancik et al., 2020; Barron et al., 2022; Martel et al., 2021; Müller et al., 2022), while $\mathcal{M}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{1}$ acts as the neural representation. Typically, $\mathcal{M}$ is an MLP.
@@
A straightforward way to solve Equation 1 is by optimizing the $J$ nodes separately, akin to individual reconstruction. However, we propose introducing a dynamic prior, which links all the models $\bm{w}_{1:J}$ during training and updates in response to their optimization, can provide a more principled approach. This Bayesian framework is designed to exploit shared statistical regularities among different objects, thereby enhancing the quality of the joint reconstruction.
@@ 3.1 Existing Methods Available for Joint Reconstruction
Let $\mathcal{G}_{\bm{\phi}}$ represent the neural representation for the static component and $\mathcal{H}_{\bm{w}}$ signify the transient component. For a given set of $J$ objects, each object-associated reconstruction node has its distinct transient network $\bm{w}_{j}$ and corresponding transient feature $\bm{b}_{j}$. In contrast, the static network $\bm{\phi}$ is shared across all nodes. The objective for this framework is formulated as:
@@
Here, $\mathcal{G}_{\bm{\phi}}^{r}(\bm{C})$ represents the static intensity, and $\mathcal{G}_{\bm{\phi}}^{\backslash r}(\bm{C})$ serves as intermediate features for the transient network. Detailed explanation and framework schematic are in Section B.2. At its core, INRWild emphasizes training the static network $\bm{\phi}$, which embodies most learnable parameters, using aggregated losses. Concurrently, the individual representations, characterized by $\bm{w}_{1:J}$ and $\bm{b}_{1:J}$, are refined based on $\bm{\phi}$’s characteristics.
@@
$\bm{\theta}^{t}=\bm{\theta}^{t-1}-\alpha\frac{1}{J}\sum_{j=1}^{J}\nabla_{\bm{w}_{j}}\ell_{j}(\bm{w}_{j}^{(K)})\frac{\partial\bm{w}_{j}^{(K)}}{\partial\bm{\theta}},$ (4)
where $\alpha$ is the outer-loop learning rate. Equation 4 is a form that can be further specialized. After $T$ steps of outer-loop optimization, the meta-learned neural representation $\bm{\theta}^{T}$ serves as an effective initialization for individual reconstructions.
@@ 4 A Novel Bayesian Framework for Joint Reconstruction
Motivation. A method that uses a composition of static and transient components operates under the assumption that all the representations substantially overlap. This may be true in 3D scene reconstruction, where observations are taken from different viewpoints of the same object. Our empirical findings on INRWild indicate such methods do not work efficiently in CT reconstruction and other image-level reconstruction tasks (discussed in Section 5). Meta-learned initialization methods train a meta model to capture a high-level representation, which can then be promptly adapted to individual objects. Although the meta model can effectively extract statistical regularities among multiple objects, we observe that this prior information tends to be lost during the adaptation phase. This loss occurs as the adaptation relies solely on local measurements, leading to an overfitting issue, as we will discuss in Section 5. Overfitting is a notorious problem in iterative methods of CT reconstruction (Herman & Odhner, 1991; Effland et al., 2020), which renders these methods sensitive to hyperparameters, such as early stopping. This sensitivity consequently increases the difficulty of deployment.
@@
Definition and Notation. We introduce distributions to the networks $\bm{w}_{1:J}$ for $J$ objects, and define latent variables $\{\bm{\omega},\bm{\sigma}\}$ that parameterize an axis-aligned multivariate Gaussian prior $\mathcal{N}(\bm{\omega},\bm{\sigma})$ from which the weights $\bm{w}_{1:J}$ are generated. A Gaussian distribution is a natural and effective choice for the weight distribution of the MLP such as INR networks (de G. Matthews et al., 2018). The latent variables collectively serve to capture the shared trends within the networks, effectively quantifying the mutual information across different objects. To simplify the model, we assume the conditional independence among all objects: $p(\bm{w}_{1:J}|\bm{\omega},\bm{\sigma})=\prod_{j=1}^{J}p(\bm{w}_{j}|\bm{\omega},\bm{\sigma})$. This assumption of conditional independence allows us to decompose the variational inference into separable optimization problems, thereby facilitating more efficient parallel computing.
Given that the measurements of the objects $\bm{y}_{1},\ldots,\bm{y}_{J}$ are mutually independent and that each network focuses on a specific object, the posterior distribution of network weights and latent variables can be derived using the Bayes’ rule as $p(\bm{w}_{1:J},\bm{\omega},\bm{\sigma}|\bm{y}_{1:J})\propto p(\bm{\omega},\bm{\sigma})\prod_{j=1}^{J}p(\bm{y}_{j}|\bm{w}_{j})p(\bm{w}_{j}|\bm{\omega},\bm{\sigma})$. While this posterior enables various forms of deductive reasoning, inferring the true posterior is often computationally challenging. Moreover, the selection of an appropriate prior $p(\bm{\omega},\bm{\sigma})$ poses its own difficulties (Wenzel et al., 2020; Fortuin et al., 2022). To tackle these issues, we present an algorithm that aims at maximizing the marginal likelihood $p(\bm{y}_{1:J}|\bm{\omega},\bm{\sigma})$, the details of derivations are provided in Appendix A.

### v2

@@ Abstract
Computed Tomography (CT) is pivotal in industrial quality control and medical diagnostics. Sparse-view CT, offering reduced ionizing radiation, faces challenges due to its under-sampled nature, leading to ill-posed reconstruction problems. Recent advancements in Implicit Neural Representations (INRs) have shown promise in addressing sparse-view CT reconstruction. Recognizing that CT often involves scanning similar subjects, we propose a novel approach to improve reconstruction quality through joint reconstruction of multiple objects using INRs. This approach can potentially utilize the advantages of INRs and the common patterns observed across different objects. While current INR joint reconstruction techniques primarily focus on speeding up the learning process, they are not specifically tailored to enhance the final reconstruction quality. To address this gap, we introduce a novel INR-based Bayesian framework integrating latent variables to capture the common patterns across multiple objects under joint reconstruction. The common patterns then assist in the reconstruction of each object via latent variables, thereby improving the individual reconstruction. Extensive experiments demonstrate that our method achieves higher reconstruction quality with sparse views and remains robust to noise in the measurements as indicated by common numerical metrics. The obtained latent variables can also serve as network initialization for the new object and speed up the learning process. [^1] We have used ChatGPT provided by OpenAI to assist in writing. The language model was employed at the sentence level for tasks such as fixing grammar and rewording sentences. We assure that all ideas, claims, and results presented in this work are human-sourced.
@@ 1 Introduction
Computed Tomography (CT) is a crucial non-invasive imaging technique, extensively employed in medical diagnostics and industrial quality control. In CT, an object’s internal structure is reconstructed from X-ray projections captured at multiple angles, posing a complicated inverse problem. In certain situations, reducing the number of CT measurements can yield advantages such as decreased radiation exposure and enhanced production throughput. However, this sparsity in measurements, along with other factors such as the presence of noise and the size of detectors, complicates the reconstruction process, making it an ill-posed inverse problem. Such challenges arise not only in CT reconstruction but also across diverse computational tasks. Therefore, while our study centers on sparse-view CT reconstruction, the core ideas are transferable to numerous inverse problems, such as Magnetic Resonance Imaging and Ultrasound Imaging.
@@
Different strategies have been developed to address the challenges of sparse-view CT reconstruction by incorporating auxiliary information. Supervised learning techniques learn mappings from measurements to images (Zhang et al., 2018b; Han & Ye, 2018; Zhu et al., 2018; Wu et al., 2021). These methods train on extensive datasets containing dense measurements, synthesizing the sparse view through physical models of the measurement process. Alternatively, Song et al. (2022) introduces an approach that directly learns the image distribution from high-quality reconstructed images using generative models. Several diffusion-based approaches (Chung et al., 2022; Liu et al., 2023; Song et al., 2023; Xu et al., 2024) have been specifically tailored for CT applications, further refining the foundational ideas presented by Song et al. (2022). However, their method still relies on a large, domain-specific dataset, that closely matches the target of reconstruction. Dependence on large datasets introduces practical limitations to supervised learning techniques and generative methods, particularly when collecting large-scale and high-quality data is intractable, such as due to privacy regulation in medical cases or timeliness in industrial settings. It is important to note that supervised learning and diffusion-based approaches represent only a subset of the methods available for CT reconstruction. There are also some studies that leverage heuristic image priors, e.g. Total Variation (TV) (Sidky & Pan, 2008; Liu et al., 2013; Zang et al., 2018). While effective, these methods lack the ability to incorporate domain-specific enhancement. Another approach involves using a few dense-view images as priors for reconstruction (Chen et al., 2008; Shen et al., 2022). However, this method hinges on the availability of dense-view images similar to the target object, which is often impractical when the appearance of the object to be reconstructed is unknown.
Building on the foundations of CT reconstruction, many works explore the potential of implicit neural representations (INRs). Thanks to the advantages of INRs, such as their continuous representation nature in contrast to conventional discrete representations (Lee et al., 2021; Xu et al., 2022; Grattarola & Vandergheynst, 2022), these methods have consistently delivered promising reconstruction results with limited number of measurements (Zang et al., 2021; Zha et al., 2022; Rückert et al., 2022; Wu et al., 2023b). Given INRs’ proven capabilities in CT reconstruction and the known advantages of incorporating prior information, either heuristically through TV (Zang et al., 2021) or explicitly from dense-view CT images (Shen et al., 2022), we seek to merge these two paradigms. Aiming for broad applicability, we avoid reliance on pre-curated datasets. Instead, we draw inspiration from the fact that modern CT machines routinely scan similar subjects, such as patients in hospitals or analogous industrial products. This commonality drives us to explore a novel question in this research:
@@
Figure 1 contrasts the methodologies of individual reconstruction against joint reconstruction. In our exploration of this research avenue, we found that several existing methods can be adapted for our purpose (Zhang et al., 2013; Ye et al., 2019; Tancik et al., 2021; Martin-Brualla et al., 2021; Kundu et al., 2022). These approaches have utilized the inherent statistical regularities across different objects, represented by the network weights in INRs. However, they often focus on addressing issues like convergence rate (Tancik et al., 2021; Kundu et al., 2022) rather than optimizing the reconstruction quality itself. As demonstrated in our experimental evaluations Section 5, these methods achieve lower PSNR and SSIM compared to our proposed method in joint reconstruction scenarios.
@@
Our Contributions: i) We explore a novel problem of INR-based joint reconstruction in the context of CT imaging, supported by a comprehensive review of existing methods that could be adapted to address this challenge. ii) We propose a principled Bayesian framework that adaptively integrates prior information throughout the training process for enhanced reconstruction quality. iii) Through extensive experiments, we evaluate various facets of reconstruction performance using common numerical metrics. Our results establish that our method outperforms the compared INR-based baselines.
@@ 2 Related Work
We briefly outline key studies related to our focal areas, with a comprehensive understanding of INR and Neural Radiance Fields (NeRF) available in the survey (Tewari et al., 2022; Molaei et al., 2023). The key working principles of INR for CT reconstruction is explained in Section 3.
Coordinate-based Multi-Layer Perceptrons (MLPs) intake spatial coordinates and output values, such as RGB or density, processed through MLP. Unlike traditional discrete representations that directly map inputs to outputs on pixel grids, these coordinate-based outputs are generated by the MLP. Therefore, such coordinate-based representations are often regarded as INRs. Originally, INRs faced challenges in capturing high-frequency details. However, the introduction of coordinate encoding, such as Fourier feature transformations (Jacot et al., 2018; Tancik et al., 2020), has significantly enhanced their ability to represent finer details.
NeRF is a state-of-the-art INR approach that reconstructs scenes from photos taken at multiple angles (Mildenhall et al., 2021; Barron et al., 2021; 2022), similar to CT where projections are acquired at different angles to compute reconstruction. While NeRF processes optical images to render RGB values by considering light transmittance and occlusion, CT reconstruction calculates attenuation coefficients to reveal internal compositions using penetrating X-rays. We therefore briefly introduce the NeRF methods that we adapt as comparison baselines. Building on the groundwork, several techniques have emerged to leverage the shared information across different objects or scenarios. NeRF-wild (Martin-Brualla et al., 2021) differentiated between static and transient scene aspects, an approach echoed in video representations (Li et al., 2021; Mai & Liu, 2022). Tancik et al. (2021) introduced meta-learning techniques to learn a meta initialization which converges faster than training from scratch. They have also applied their method to CT reconstruction. For the same purpose of accelerating convergence, Kundu et al. (2022) introduced federated learning to obtain the meta initialization. Chen & Wang (2022) proposed using transformers as the meta-learners.
@@ 3 Problem Statement and Preliminaries
The CT acquisition process, in an ideal noiseless scenario, can be mathematically described by the equation: $\bm{y}=\bm{A}\bm{x}$, where $\bm{x}\in{\mathbb{R}}^{m}$ represents the unknown object of interest and $\bm{y}\in{\mathbb{R}}^{n}$ symbolizes the ideal noiseless measurements. In practice, the actual obtained measurements $\widetilde{\bm{y}}\in{\mathbb{R}}^{n}$ are often noisy, where the difference between actual and ideal measurements $\bm{\epsilon}=\widetilde{\bm{y}}-\bm{y}$ is the measurement noise. We note that the measurement noise $\bm{\epsilon}\in{\mathbb{R}}^{n}$ is a stochastic term representing general noise, including non-additive types like Poisson noise, which is the primary focus of this study. For simplicity and to maintain focus on the methodological aspects, this manuscript primarily expands formulas under the assumption of noiseless measurements $\bm{y}$, unless specified otherwise.
These measurements arise from the interaction between the measurement matrix $\bm{A}\in{\mathbb{R}}^{n\times m}$ and the object. The task in CT is to infer the unknown object $\bm{x}$ from the acquired CT measurements $\bm{y}$. The inherent challenge lies in the common sparsity of these measurements, resulting in $m>n$. This makes the reconstruction problem ill-posed. Meanwhile, measurement noise can further corrupt the reconstruction results.
The INR designed for CT reconstruction is a function $f_{\bm{w}}:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}^{1}$ parameterized by $\bm{w}$. It maps the three-dimensional spatial coordinates of the object to its intensity. INR consists of two components, formulated as $f_{\bm{w}}=\mathcal{M}\circ\gamma$. Here, $\gamma:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}^{d}$ servers as the position encoding (Tancik et al., 2020; Barron et al., 2022; Martel et al., 2021; Müller et al., 2022), while $\mathcal{M}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{1}$ acts as the neural representation. Typically, $\mathcal{M}$ is an MLP.
@@
While this formulation allows for different measurement matrices for each object, we primarily consider scenarios where all objects share the same measurement matrix $\bm{A}$ in our experiments. A straightforward way to solve Equation 1 is by optimizing the $J$ nodes separately, akin to individual reconstruction. However, we propose introducing a dynamic prior, which links all the models $\bm{w}_{1:J}$ during training and updates in response to their optimization. This Bayesian framework is designed to exploit shared statistical regularities among different objects, thereby enhancing the quality of the joint reconstruction.
@@ 3.1 Existing Methods Available for Joint Reconstruction
Let $\mathcal{G}_{\bm{\phi}}$ represent the neural representation for the static component ($\bm{\phi}$ is the weights of MLP $\mathcal{G}$) and $\mathcal{H}_{\bm{w}}$ signify the transient component ($\bm{w}$ is the weights of MLP $\mathcal{H}$). For a given set of $J$ objects, each object-associated reconstruction node has its distinct transient network $\bm{w}_{j}$ and corresponding transient feature $\bm{b}_{j}$. In contrast, the static network $\bm{\phi}$ is shared across all nodes. The objective for this framework is formulated as:
@@
The output of the static network $\mathcal{G}_{\bm{\phi}}$ is split into two components: $\mathcal{G}_{\bm{\phi}}^{r}(\bm{C})$, which represents the static intensity, and $\mathcal{G}_{\bm{\phi}}^{\backslash r}(\bm{C})$, which serves as intermediate features for the transient networks. The symbols $r$ and \$\backslash r$ denote the division of outputs into these two components. Detailed explanation and framework schematic are in Section B.2. At its core, INRWild emphasizes training the static network $\bm{\phi}$, which embodies most learnable parameters, using aggregated losses. Concurrently, the individual representations, characterized by $\bm{w}_{1:J}$ and $\bm{b}_{1:J}$, are refined based on $\bm{\phi}$ that corresponds to the static part. $\bm{w}_{1:J}$, $\bm{b}_{1:J}$ and $\bm{\phi}$ are jointly optmized to achieve the composition of static and transient parts.
@@
$\bm{\theta}^{(t)}=\bm{\theta}^{(t-1)}-\alpha\frac{1}{J}\sum_{j=1}^{J}(\nabla_{\bm{w}_{j}})^{T}\ell_{j}(\bm{w}_{j}^{(K)})\frac{\partial\bm{w}_{j}^{(K)}}{\partial\bm{\theta}},$ (4)
where $\alpha$ is the outer-loop learning rate. Equation 4 is a form that can be further specialized. After $T$ steps of outer-loop optimization, the meta-learned neural representation $\bm{\theta}^{(T)}$ serves as an effective initialization for individual reconstructions.
@@ 4 A Novel Bayesian Framework for Joint Reconstruction
Motivation. NeRFWild is proposed to address scene reconstruction challenges, such as reconstructing popular sightseeing sites from in-the-wild photos. These scenarios often involve a primary target object, like a landmark, amidst various transient elements such as pedestrians and changing light conditions. A method that uses a composition of static and transient components operates under the assumption that the core representations of the scene—those of the main object—substantially overlap, despite different viewpoints or transient changes. This assumption typically holds true in 3D scene reconstruction, where observations are collected from multiple perspectives of the same object. Our empirical findings on INRWild indicate such methods do not work efficiently in CT reconstruction and other image-level reconstruction tasks (discussed in Section 5). Meta-learned initialization methods, such as FedAvg and MAML, train a meta model to capture a high-level representation, which can then be promptly adapted to individual objects. Although the meta model can effectively extract statistical regularities among multiple objects, we observe that this prior information tends to be lost during the adaptation phase. This loss occurs as the adaptation relies solely on local measurements, leading to an overfitting issue, as we will discuss in Section 5. Overfitting is a notorious problem in iterative methods of CT reconstruction (Herman & Odhner, 1991; Effland et al., 2020), which renders these methods sensitive to hyperparameters, such as early stopping. This sensitivity consequently increases the difficulty of deployment.
@@
Definition and Notation. We introduce distributions to the networks $\bm{w}_{1:J}$ for $J$ objects, and define latent variables $\{\bm{\omega},\bm{\sigma}\}$ that parameterize an axis-aligned multivariate Gaussian prior $\mathcal{N}(\bm{\omega},\bm{\sigma})$ from which the weights $\bm{w}_{1:J}$ are generated. Each node is assumed to be of the same size. A Gaussian distribution is a natural and effective choice for the weight distribution of the MLP such as INR networks (de G. Matthews et al., 2018). The latent variables collectively serve to capture the shared trends within the networks, effectively quantifying the mutual information across different objects. To simplify the model, we assume the conditional independence among all objects: $p(\bm{w}_{1:J}|\bm{\omega},\bm{\sigma})=\prod_{j=1}^{J}p(\bm{w}_{j}|\bm{\omega},\bm{\sigma})$. This assumption of conditional independence allows us to decompose the variational inference into separable optimization problems, thereby facilitating more efficient parallel computing.
Given that the measurements of the objects $\bm{y}_{1},\ldots,\bm{y}_{J}$ are mutually independent and that each network focuses on a specific object, the posterior distribution of network weights and latent variables can be derived using the Bayes’ rule as $p(\bm{w}_{1:J},\bm{\omega},\bm{\sigma}|\bm{y}_{1:J})\propto p(\bm{\omega},\bm{\sigma})\prod_{j=1}^{J}p(\bm{y}_{j}|\bm{w}_{j})p(\bm{w}_{j}|\bm{\omega},\bm{\sigma})$. While this posterior enables various forms of deductive reasoning such as reconstruction uncertainty quantification, inferring the true posterior is often computationally challenging. Moreover, the selection of an appropriate prior $p(\bm{\omega},\bm{\sigma})$ poses its own difficulties (Wenzel et al., 2020; Fortuin et al., 2022). To tackle these issues, we present an algorithm that aims at maximizing the marginal likelihood $p(\bm{y}_{1:J}|\bm{\omega},\bm{\sigma})$, integrating out $\bm{\omega}_{1:J}$. Detailed derivations are provided in Appendix A.